Operation Summary for January 24, 2011 - January 26, 2011
Monday Jan. 24
The day shift began with Operations monitoring the Tevatron (TEV) store 8443 (54.78E30) and the Recycler stash (345.0E10), with the Antiproton Source (Pbar) stacking (9.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNEexperiments taking beam, and Meson FTBF T-978 (CALICE)experiment off due to beamline vacuum problems, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the day shift, Operations terminated store 8443 at 9:17 AM, put the TEV into HEP shot setup, and established store 8444 at 10:25 AM with an average initial luminosity (AIL) of 328.41E30. Linac personnel switched over from the I- Source to the H- Source at 11 AM. Water Group personnel completed repairs on the NuMI LCW expansion tank at 11:44 AM.During the evening shift, Operations monitored the accelerator complex.
Tuesday Jan. 25
The midnight shift began with Operations monitoring store 8444 (85.08E30) and the stash (347.5E10), with Pbar stacking (17.5E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8444 at 1:45 AM, put the TEV into shot setup, and established store 8445 at 3:02 AM with an AIL of 339.93E30.
Wednesday Jan. 26
The midnight shift began with Operations monitoring store 8446 (185.58E30) and the stash (122.9E10), with Pbar stacking (19.0E10), and with NuMI and MiniBooNE taking beam.During the midnight shift, Operations monitored the accelerator complex.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 01/26/2011 email Fermilab|
|Security, Privacy, Legal|